OpenStudy (anonymous):

differentiate: a^4y=(x^2+a^2)^3/2

hartnn (hartnn):

diff w.r.t y, a or x ?

OpenStudy (anonymous):

differentiate y

hartnn (hartnn):

ok,assuming diff y w.r.t x first step is to take log on both sides.

OpenStudy (anonymous):

how ?

hartnn (hartnn):

\[4y*\log a=\frac{3}{2}\log(x^2+a^2)\]

OpenStudy (anonymous):

why is it log?

hartnn (hartnn):

is the question,why is it not ln ? or why did i take log altogether?

OpenStudy (anonymous):

the second one .. ahm, is the log part of differential ?

hartnn (hartnn):

actually for the 1st part,it should be ln and not log,to indicate logarithm to base e i wanted to isolate y,thats y i took log on both sides,as u can see now,u can write y=........... also ln is the part of differentiation \[\frac{d}{dx}\ln f(x)=\frac{1}{f(x)}\frac{d}{dx}f(x)\]

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

i don't know .. tsk . . dumb!